OpenStudy (anonymous):

A ball is dropped from the top of a 253-foot building. How many seconds will it take to hit the ground if the height is given by h=-16t^2+s where s is the initial height? answer in unit of seconds

OpenStudy (anonymous):

253 feet= 77.114 meters. and accel due to gravity = -9.8 m/s^2 final height = 0m 0m= 1/2(-9.8)t^2+0t+77.114 t=3.9607 seconds

OpenStudy (anonymous):

thanks!